1.) A gas is found to have a density of 3.165 g/L at STP. What is its molar mass? Which of the following gases is it most likely to be?

Molar Mass: _________

a. NH3
b. O2
c. Cl2

d. SF6

2.) Calculate the pressure exerted by 5.05 moles of NCl3 gas in a 20.00-L container at -205°C for a real gas. (for xenon: a= 42.63 atm·L2/mol2 and b= 0.2450 L/mol)

3.) Which of the following is NOT true in the kinetic molecular theory of ideal gases?

The size of the molecules are negligible compared to the volume of the container

The kinetic energy of a sample is proportional to the temperature of the gas.

Collisions are elastic (no friction occurs between the molecules when they collide)

The pressure of a sample is affected by which gases are present.
Answer #1

1)

P = 1.0atm

T = 273 K

density = 3.165 g/L

Lets derive the equation to be used

we have below equation to be used:

p*V=n*R*T

p*V=(mass/molar mass)*R*T

p*molar mass=(mass/V)*R*T

p*molar mass=density*R*T

Put Values:

1.0 atm * MM = density * 0.0821 atm.L/mol.K *273.0 K

MM = 70.9381 g/mol

This is molar mass of Cl2

Answer: c

Feel free to comment below if you have any doubts or if this answer do not work

Only 1 question at a time please
Answer #2

cl2 is the correct option

answered by: Shivani Sharma